A car enters a curve in the road with a speed of 44 m/sm/s and emerges from this curve 16 s later with the same speed. However, the direction of the velocity changes by 115 degrees during this time.
What is the magnitude of the average acceleration during this time interval?
Answer #1

magnitude of average aceleration,

a_avg = |vf - vi|/t

|vf - vi| = sqrt(vf^2 + vi^2 - 2*vf*vi*cos(theta))

= sqrt(44^2 + 44^2 - 2*44*44*cos(115))

= 74.22 m/s

a_avg = 74.22/16

= 4.64 m/s^2 <<<<<<<<<<<---------------------Answer